OpenStudy (anonymous):

Help please!! Identify the vertex for the graph of y = 2x^2 + 8x - 3. (2, 21) (2, 17) (-2, -11) (-2, -27)

OpenStudy (anonymous):

You can get the vertex by the formula -b/21, c-(b^2/41) For the x coordinate, -8/4 = -2 For the y coordinate, -3 - 64/8 = -3 - 8 = -11 So, (-2, -11)

OpenStudy (anonymous):

What @raechelvictoria said lol

OpenStudy (anonymous):

You can also just find the x coordinate, which is 2, then substitute it to the original equation to find for the y coordinate.

OpenStudy (anonymous):

*to find the y coordinate

OpenStudy (anonymous):

*-2

OpenStudy (anonymous):

OMG sorry for all the typos lol

OpenStudy (anonymous):

haha its ok & thank you!

OpenStudy (anonymous):

Can you help me with another one? @raechelvictoria

OpenStudy (anonymous):

Okay sure :)

OpenStudy (anonymous):

thanks :) The function f(x) = -x^2 - 9x + 10 shows the relationship between the vertical distance of a diver from a pool's surface f(x), in feet, and the horizontal distance x, in feet, of a diver from the diving board. What is a zero of f(x) and what does it represent? x = 1; the diver jumps in the pool at 1 foot per second x = 10; the diver jumps in the pool at 10 feet per second x = 1; the diver hits the water 1 foot away horizontally from the board x = 10; the diver hits the water 10 feet away horizontally from the board

OpenStudy (anonymous):

The third one.

OpenStudy (anonymous):

You have to factor the equation, which will give us (x+10)(x-1), which gives us x=-10, 1. Since there is no negative distance, we only have x = 1 ft. The problem states that x represents horizontal distance x, which is why the third choice is the answer.

OpenStudy (anonymous):

thanks!